# Generative Evolutionary Meta-Solver (GEMS): Scalable Surrogate-Free Multi-Agent Learning

### v1

@@ 5 Conclusion
Our approach removes the linear memory growth and quadratic computation overhead inherent to classical Psro while preserving key game-theoretic guarantees. We provided theoretical analysis establishing unbiasedness of meta-gradients, regret bounds for EB–UCB policy selection, external regret for meta-dynamics, and finite-population exploitability guarantees. Empirically, Gems demonstrates faster convergence, lower memory footprint, and improved scalability across challenging multi-agent benchmarks.
@@
Acknowledgement
The authors also wish to acknowledge the use of ChatGPT/Gemini in the writing of this paper. This tool was used to improve and polish the presentation of text, tables and figures in the paper. The paper remains an accurate representation of the authors’ underlying work and novel intellectual contributions.
Ethics Statement
Our research focuses on improving the efficiency and scalability of multi-agent reinforcement learning in simulated environments, without involving human subjects or sensitive personal data. All experiments are conducted in controlled, synthetic settings, and we do not deploy agents in real-world competitive domains. We prioritize transparency and reproducibility and will be providing full code, models, and hyperparameters, enabling others to validate and build upon our results. Additionally, by reducing memory and computation overhead compared to classical PSRO, our method promotes more resource-efficient research. We emphasize that responsible application of multi-agent learning frameworks is essential, and any use in strategic or competitive real-world systems should carefully consider fairness, safety, and potential misuse.
Reproducibility Statement
All code, models, and data required to reproduce our experiments will be publicly released upon acceptance. Our implementation uses Python 3.11.9 and PyTorch 2.8, with all dependencies listed in requirements.txt. We will provide full details of hyperparameters, training schedules, rollout counts, and random seeds, ensuring that all reported results, including memory efficiency, convergence, and exploitability metrics, can be reproduced. Evaluation scripts for Monte Carlo payoff estimation and tournament analyses will be included to allow full replication of the figures and tables in the paper.
Appendix
Appendix - Table of Contents
@@ Acknowledgement
This appendix provides the detailed mathematical analysis supporting the claims made in Section 3. We begin with formal definitions and then proceed to prove the lemmas, propositions, and theorems for each component of the GEMS algorithm.
A Formal Preliminaries
@@ A.2 Nash Equilibrium and Exploitability
B Proofs for Meta-Game Estimation
@@ B.1 Supporting Definitions: Empirical-Bernstein Inequality
C Proofs for Meta-Game Solving
@@
The standard OMWU analysis (e.g., in Hazan, "Introduction to Online Convex Optimization" ([2])) shows that these steps lead to a bound on the regret against the estimated losses:
@@
D Proofs for the Bandit Oracle
@@
E Proof of Overall Exploitability
@@
F Extension to Two-Player General-Sum Games
@@ F.4 Overall Guarantee: $\epsilon$-Coarse-Correlated Equilibrium
G Proofs for Two-Player General-Sum
@@ G.4 Proof of Theorem 4 ($\epsilon$-CCE)
H Extension to N-Player General-Sum Games
@@ H.3 Overall Guarantee: $\epsilon$-CCE for n Players
I Proofs for N-Player General-Sum
@@ I.4 Proof of Theorem 6 ($\epsilon$-CCE for n Players)
Part III Experiments
J Coordination on Simple Spread

### v2

@@ 5 Conclusion
Our approach removes the linear memory growth and quadratic computation overhead inherent to classical Psro while preserving key game-theoretic guarantees. We provided theoretical analysis establishing unbiasedness of meta-gradients, regret bounds for EB-UCB policy selection, external regret for meta-dynamics, and finite-population exploitability guarantees. Empirically, Gems demonstrates faster convergence, lower memory footprint, and improved scalability across challenging multi-agent benchmarks.
@@
6 Limitations
While Gems presents a significant step toward scalable, surrogate-free multi-agent learning, we acknowledge the limitations that define the scope of this work and offer avenues for future research.
  • Benchmark Scope and Baselines: Our empirical evaluation focuses on demonstrating the fundamental scalability and game-solving advantages of Gems over the Psro paradigm. Consequently, some of our experiments (e.g., Multi-Agent Tag and Simple Spread) use classical PSRO as the primary baseline. This choice is deliberate, as classical Psro perfectly embodies the core $O(k^{2})$ computational and $O(k)$ memory bottlenecks that Gems is designed to solve. To validate GEMS’s strategic performance, we also benchmarked it against a suite of modern, stronger Psro variants on complex games like Kuhn Poker and the Deceptive Messages Game. However, we did not conduct experiments on large-scale benchmarks such as the StarCraft Multi-Agent Challenge (SMAC). The prohibitive computational cost of training even a single baseline Psro agent on such complex maps made a direct comparison infeasible with our available resources.
  • Hyperparameter Sensitivity: Gems introduces a new set of hyperparameters related to the amortized generator, the ABR-TR objective (e.g., $\beta$, $\mathcal{L}_{\text{ABR-TR}}$ learning rate), and the EB-UCB oracle (e.g., candidate pool size $|\Lambda_{t}|$). While we provide ablation studies in the Appendix to analyze the sensitivity of key parameters in environments like Kuhn Poker and the Public Goods Game, a comprehensive, large-scale analysis of the interplay between all hyperparameters across diverse game types is beyond the scope of this initial work.
  • Generator Capacity: The efficacy of Gems relies on the capacity of the single amortized generator $G_{\theta}$ to represent a rich, continuous space of diverse and effective policies. This work utilized standard MLP architectures for the generator. The performance of Gems in much more complex domains (e.g., with high-dimensional observation/action spaces) may depend heavily on the choice of more advanced generator architectures (e.g., Transformers, diffusion models). Exploring the architectural limits of the generator is a key direction for future work.
  • Candidate Pool Generation: The EB-UCB oracle’s ability to find effective new strategies is dependent on the quality of the candidate pool $\Lambda_{t}$ it searches over. Our current implementation relies on simple heuristics for generating this pool (e.g., mutation and random sampling, as detailed in the Appendix ablations). The performance of Gems could potentially be improved by incorporating more sophisticated or guided methods for generating candidate latent codes.
Appendix – Table of Contents
@@ 6 Limitations
This appendix provides the detailed mathematical analysis supporting the claims made in §3. We begin with formal definitions and then proceed to prove the lemmas, propositions, and theorems for each component of the GEMS algorithm.
Appendix A Formal Preliminaries
@@ A.2 Nash Equilibrium and Exploitability
Appendix B Proofs for Meta-Game Estimation
@@ B.1 Supporting Definitions: Empirical-Bernstein Inequality
Appendix C Proofs for Meta-Game Solving
@@
The standard OMWU analysis (e.g., in Hazan, "Introduction to Online Convex Optimization" (Hazan (2023))) shows that these steps lead to a bound on the regret against the estimated losses:
@@
Appendix D Proofs for the Bandit Oracle
@@
Appendix E Proof of Overall Exploitability
@@
Appendix F Extension to Two-Player General-Sum Games
@@ F.4 Overall Guarantee: $\epsilon$-Coarse-Correlated Equilibrium
Appendix G Proofs for Two-Player General-Sum
@@ G.4 Proof of Theorem 4 ($\epsilon$-CCE)
Appendix H Extension to N-Player General-Sum Games
@@ H.3 Overall Guarantee: $\epsilon$-CCE for n Players
Appendix I Proofs for N-Player General-Sum
@@ I.4 Proof of Theorem 6 ($\epsilon$-CCE for n Players)
Part III Ablation and Analysis of experiments
All empirical evaluations presented in this paper were conducted on computing systems equipped with Intel i9 processors, utilizing either NVIDIA RTX A2000 GPU cluster (12GB VRAM) or an NVIDIA RTX 6000 Ada Generation cluster (48GB VRAM).
Table 2: Comprehensive Summary of Gems Experiments. This table aggregates all environments evaluated in the paper, categorized by scale and type. It confirms that Gems scales robustly to high-dimensional latent spaces ($d_{z}=16$) and large populations across competitive, cooperative, and game-theoretic ablation settings.
Environment Type Action Space Latent Dim ($d_{z}$) Iterations ($T$) Policies ($|Z_{T}|$)
Complex Scale
Chess (v6) Competitive $\approx 4,672$ 8 1,000 2,002
Go (9x9) Competitive $\approx 82$ 8 200 402
Hanabi Cooperative $\approx 20$ 8 200 402
Connect 4 Competitive 7 8 200 402
Continuous Control (Multi-particle)
Simple Spread (MPE) Cooperative Continuous (5) 8 100 202
Simple Tag (MPE) Competitive Continuous (5) 8 100 202
Game-Theoretic and Ablations
Kuhn Poker Imperfect Info 2 2, 4, 8 40 82
Public Goods Game N-Player Continuous 8, 16 10 60
Deceptive Mean Zero-Sum Continuous 8, 16 6 14
Note: "Policies ($|Z_{T}|$)" is calculated as $|Z_{0}|+N_{players}\times T$.
Initialization of the anchor set.
Algorithm 1 takes the initial anchor set $\mathcal{Z}_{0}=\{z_{1},\dots,z_{k_{0}}\}$ as an input. We intentionally leave the choice of this initialization to the implementer: anchors can be sampled from a prior over the latent space, warm-started from pretrained policies, or constructed via domain-specific heuristics. In our experiments, we use a simple and standard choice and initialize anchors independently from a standard normal distribution, i.e., $z_{i}\sim\mathcal{N}(0,I)$, with the initial meta-strategy $\sigma_{0}$ set to be uniform over $\mathcal{Z}_{0}$.
Initialization of the meta-strategy.
The initial meta-strategy $\sigma_{0}$ is defined as a probability distribution over the initial anchor set $\mathcal{Z}_{0}$. In our experiments, we initialize $\sigma_{0}$ to be uniform over $\mathcal{Z}_{0}$, assigning equal probability mass to each initial anchor. This choice reflects the absence of prior preference among anchors at initialization and provides a neutral starting point for subsequent meta-strategy updates via optimistic multiplicative weights. Alternative initializations, such as biased or warm-started distributions, can be incorporated without modifying the core algorithm.
Policy optimization in the amortized best-response.
In Phase 4 of Algorithm 1, agents are trained via policy-gradient methods using advantage-based objectives. Concretely, the generator parameters $\theta$ are updated by ascending a regularized policy objective that depends on estimated advantages, allowing any standard policy-based reinforcement learning algorithm to be used in this step. In our implementation, we instantiate this update using Proximal Policy Optimization (PPO) with generalized advantage estimation (GAE), and train agents in PettingZoo environments accordingly. We emphasize that this choice is not intrinsic to Gems, and alternative policy-gradient or trust-region methods could be substituted without altering the overall framework.
Default hyperparameters for Kuhn Poker.
For reproducibility, Table 3 reports the exact default hyperparameters used in our Kuhn Poker runs of Gems (Algorithm 1). Unless otherwise stated, all Kuhn experiments in this paper use these settings: an outer-loop budget of T=40T{=}40 with $k_{0}{=}1$ initial anchor per role (implying $|Z_{T}|{=}2(k_{0}{+}T){=}82$ anchors for the two-player case), latent dimension $d_{z}{=}8$, and fixed Monte Carlo / oracle budgets for meta-game estimation and EB-UCB expansion. Meta-strategy updates use OMWU with $\eta{=}0.03$ and a constant schedule, while Phase 4 applies ABR-TR training for $K_{\text{ABR}}{=}30$ gradient steps per outer iteration with learning rate $2\times 10^{-4}$ and KL regularization weight $\beta_{\text{KL}}{=}0.05$ (Table 3).
Table 3: Default hyperparameters for GEMS (Algorithm 1) on Kuhn poker. Defaults used in the reference implementation, grouped by phases: (1) MC meta-game estimation, (2) OMWU update, (3) EB-UCB expansion, (4) ABR-TR training.
Block Parameter Symbol Default Role
Budget / population
Iterations Outer iterations $T$ 40 Outer-loop horizon.
Initial anchors (per role) Init. anchors $k_{0}$ 1 Initial anchor set size.
Max anchors (per role) Cap $k_{\max}$ 32 Caps population size.
Total anchors (2-player) Implied total $|Z_{T}|$ $2(k_{0}+T)=82$ $|Z_{T}|=|Z_{0}|+2T$, $|Z_{0}|=2k_{0}$.
Generator / latent
Latent dimension Latent size $d_{z}$ 8 Latent code dimension.
Temperature Temp. $\tau$ 1.0 Sampling temperature.
Phase 1: Meta-game estimation (MC)
Opponent batch size Opp. batch $N_{\text{MC}}$ 8 Opponents from $\sigma_{t-1}$.
Rollouts per matchup Rollouts $M_{\text{MC}}$ 2 Rollouts per matchup.
Bootstrap samples Bootstraps $B$ 128 Bootstrap resamples.
EMA smoothing EMA coef. $\alpha_{\text{EMA}}$ 0.0 0 disables EMA.
Phase 2: Meta-strategy update (OMWU)
OMWU step size Step size $\eta$ 0.03 OMWU learning rate.
Schedule Schedule const const/sqrt/harmonic.
Logit cap Logit clip 50.0 Caps logits.
Phase 3: Expansion (EB-UCB)
Oracle opponent batch size Opp. batch $N_{\text{oracle}}$ 8 Opponents from $\sigma_{t}$.
Oracle rollouts per candidate Rollouts $M_{\text{oracle}}$ 2 Rollouts per candidate.
Mutated candidates Mut. pool $|\Lambda_{\text{mut}}|$ 32 Mutation candidates.
Random candidates Rand. pool $|\Lambda_{\text{rand}}|$ 32 Extra random candidates.
Mutation scale Mut. stddev $\sigma_{\text{mut}}$ 0.2 Latent mutation radius.
UCB confidence Confidence $\delta_{0}$ 0.5 UCB confidence.
Phase 4: ABR-TR training
Training steps Grad steps $K_{\text{ABR}}$ 30 Updates per outer iter.
Anchors per batch Batch anchors 16 Anchors from $Z_{t}$.
Learning rate Optimizer lr $2\times 10^{-4}$ Step size for generator.
KL penalty weight KL coef. $\beta_{\text{KL}}$ 0.05 Trust-region weight.
New-opponent fraction Mix frac. $q_{\text{new}}$ 0.25 Mix new vs. old.
Gradient clipping Norm clip 0.5 Global grad-norm clip.
Appendix J Run on Connect–4
Additional Analysis on Connect-4 (PettingZoo).
Figures 79 present supplementary diagnostics for Gems on the Connect-4 environment implemented using PettingZoo (connect_four_v3), evaluated across four seeds (0–3). Figure 7 compares memory consumption and cumulative wall-clock time over training iterations for Gems and Psro. While Psro exhibits steadily increasing memory usage and superlinear growth in cumulative runtime—a consequence of explicit population expansion, payoff table maintenance, and repeated best-response computation—Gems maintains near-constant memory throughout training and achieves substantially lower cumulative wall-clock time. This behavior is consistent with the amortized nature of Gems, where a single generator implicitly represents and updates the evolving policy population without requiring explicit storage or enumeration of all past strategies.
Figures 8 and 9 visualize principal component projections of learned representations across seeds, providing insight into the relationship between behavioral convergence and representational structure. In Figure 8, the PCA embeddings of action-distribution features reveal a continuous, non-linear manifold structure rather than disjoint clusters. Since Connect-4 is a solved game (perfect information, zero-sum) where optimal play dictates specific minimax trajectories, the observed strong overlap between players is theoretically expected: it indicates that Gems has converged to the shared, symmetric geometry of optimal play. The smooth gradients visible in the embeddings suggest that the generator $\mathcal{G}_{\theta}$ does not merely memorize a single solution point, but captures the smooth topology of valid strategies leading to the Nash equilibrium.
In contrast, Figure 9 shows that the corresponding latent-space embeddings remain more dispersed across seeds and players. This increased dispersion indicates that, while the induced action distributions converge to the singular optimal behavior required by the solved game mechanics, the underlying latent representations preserve diversity. This prevents the generator from collapsing into a narrow region of the latent space, ensuring it retains the expressivity to generate counter-strategies if the opponent were to deviate from optimality.
Taken together, these results illustrate a key property of Gems: Behavioral convergence emerges at the level of action distributions (reflecting the solved nature of the game), while representational diversity is retained within the latent space. This decoupling enables Gems to scale efficiently without sacrificing expressivity, supporting stable learning dynamics and reducing the risk of premature mode collapse. Combined with the resource efficiency observed in Figure 7, these findings reinforce the advantage of amortized, generator-based meta-solvers over classical Psro-style approaches in multi-agent zero-sum games.
Figure 7: Connect–4: Resource Usage. Memory consumption (left) and cumulative wall-clock time (right) over training iterations for Gems and Psro.
Connect–4: Action Embeddings (Seeds 0–3)
Figure 8: Connect–4: Learned Strategy Manifold. PCA visualization of state-conditional action distributions across four random seeds (0–3). Unlike discrete population methods, the embeddings reveal a continuous, non-linear manifold structure, indicating that the generator $\mathcal{G}_{\theta}$ captures the complex geometry of the solution space. The smooth gradients in the projection correspond to coherent variations in strategic aggression and confidence, confirming that Gems maintains structured diversity and avoids mode collapse while converging to symmetric equilibrium behaviors.
Connect–4: Latent Embeddings (Seeds 0–3)
Figure 9: Connect–4: Latent-Space Embeddings. PCA visualization of learned latent representations across four seeds (0–3). Compared to action space, the latent space remains more dispersed, suggesting preserved representational diversity despite similar cooperative behavior.
Appendix K Run on Hanabi
Additional Analysis on Hanabi (PettingZoo).
Figures 1012 present supplementary diagnostics for Gems on the cooperative Hanabi environment implemented using PettingZoo, evaluated across four seeds (0–3). Figure 10 reports memory consumption and cumulative wall-clock time over training iterations for Gems and Psro. As in the Connect-4 setting, Psro exhibits steadily increasing memory usage and superlinear growth in cumulative runtime, reflecting the cost of explicit population expansion, policy storage, and repeated equilibrium computation. In contrast, Gems maintains near-constant memory throughout training and achieves substantially lower cumulative wall-clock time, despite the increased complexity of Hanabi arising from partial observability, information asymmetry, and larger joint action spaces. These results highlight the scalability benefits of Gems in cooperative multi-agent environments where classical Psro-style methods incur significant computational overhead.
Figures 11 and 12 further analyze the learned representations through PCA visualizations of action-distribution features and latent embeddings, respectively. In contrast to perfect-information games like Connect-4 where symmetric self-play leads to overlapping embeddings, Figure 11 reveals that the action-distribution embeddings for the two players occupy distinct, complementary regions of the manifold. This separation is geometrically consistent across seeds and is expected: Hanabi is a game of asymmetric information, where each player possesses private knowledge (their partner’s cards) that strictly differentiates their optimal policy from that of their partner.
Consequently, the plots demonstrate that Gems has successfully learned a complementary coordination protocol. The smooth, symmetric gradients visible in the plots suggest the generator has discovered a continuous “convention” —a shared latent topology that maps distinct agent perspectives to compatible joint actions. This confirms that Gems avoids the mode collapse often seen in generative cooperative learning, instead maintaining the structured diversity required to solve partially observable settings where no single static policy is optimal.
Similarly, Figure 12 shows that the corresponding latent-space embeddings exhibit a marked contrast: unlike the separated action spaces, the latent codes for both players are densely intermingled. This indicates that the generator utilizes a shared latent protocol (or codebook) to encode the game state, from which distinct, complementary roles emerge in the action space. The preservation of diversity within this shared latent region suggests that Gems maintains a rich internal representation capable of supporting asymmetric coordination without fracturing into disjoint policies.
Taken together, these results demonstrate that Gems achieves efficient and stable learning in Hanabi by discovering structured, complementary manifolds. The ability to learn these implicit conventions without explicit communication channels underscores the suitability of amortized, generator-based meta-solvers for complex cooperative multi-agent environments with partial observability.
Figure 10: Hanabi: Resource Usage. Memory consumption (left) and cumulative wall-clock time (right) over training iterations for Gems and Psro.
Hanabi: Action Embeddings (Seeds 0–3)
Figure 11: Hanabi: Cooperative Coordination Manifold. PCA visualization of action-distribution features for both players across four random seeds. The embeddings reveal a coherent joint manifold with smooth entropy gradients. While the separation between players reflects the asymmetric information inherent to Hanabi, the symmetric and complementary geometry of the clusters demonstrates that the generator $\mathcal{G}_{\theta}$ captures a unified coordination protocol (or “convention”) shared by both agents, enabling zero-shot coordination despite distinct individual perspectives.
Hanabi: Latent Embeddings (Seeds 0–3)
Figure 12: Hanabi: Latent-Space Embeddings. PCA visualization of learned latent representations across four seeds (0–3). Compared to action space, the latent space remains more dispersed, suggesting preserved representational diversity despite similar cooperative behavior.
Appendix L Coordination on Simple Spread